A bank offers APR of 6.5%.
By what decimal number do you multiply a deposit to find the total amount in an account after one year at this rate?
Thanks!!!!

Answer by stanbon(75887)

A bank offers APR of 6.5%.
By what decimal number do you multiply a deposit to find the total amount in an account after one year at this rate?
--------
A(t) = P(1+r*t)
----
A(1) = P(1+0.065)
----
A(1) = P(1.065)
-----
decimal = 1065/1000
-----------------------
Cheers,
Stan H.